RC meeting and Run Plan, Friday October 28th
From SBS wiki
Return to the RC Meeting Calendar
RC Meeting Information
Time:at 3:00pm today
- Location: Counting house conference room
- Join ZoomGov Meeting
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
- Current Polarization ~31%
- Requesting hourly screenshots of Target Strip Charts (and keep an eye on these during shift)
- Work continues to optimize: updated needed today?
- Small fire took beam down ~1am; damage seen originally fixed around noon, but it was discovered that additional repairs needed. Last estimate (~12:30pm) was two more hours for the work.
- GEM tests(?) this afternoon: report?
- Other Subsystem updates?
- Redo Unser calibration? - Simona
- Luminosity scans with GEMs - Simona and Holly planning for Sunday - UPDATE: DECIDE TO DELAY AND DO THIS ONLY IF WE HAVE BEAM BUT CAN'T TAKE PRODUCTION DATA
- "Approved to do SBS GEM lumi scan using C-optics foil when the polarized target is down for say a spin up or some other reason and we cannot do production". - Simona (will/did?) look at optics foil position during beam recovery.
- (Opportunistic BCM calibration (higher current) before pass change?)
- (INFN LV Scan and DAQ Stability - maybe 20-30 minutes if nothing else can run) UPDATE: CAN BE DONE USING THE STAND-ALONE DAQ, SO NO LONGER ON THE OPPORTUNISTIC LIST.
- (Misc. GEM commissioning items: https://sbs.jlab.org/wiki/index.php/CommissioningPlan#BB.2FSBS_GEM_Commissioning_.28Continuation.29)
- Configuration and pass change on Monday (10/31/2022): 
- Significant configuration change (SBS, HCal, BB, target, plus smaller jobs) - 3 days schedule. Restoration Wednesday swing?
- Swapping of the leads to the power supply of the 3rd Moller quad
- Replace oven heater interlock box
- Plans for Monday Owl? Magnet mapping? - Arun/Gordon/??
Current Run Plan
NMR measurement Restore beam (standard post-downtime procedure)
Run Plan 1.) He-3 runs, at 45uA of beam current. Use ps3==0, ps5=12, all others -1. Take four runs that should each last for one hour. 2.) Move the target to the pickup coil position, and the TO should run an NMR sweep to measure the polarization. 3.) Return to step 1.)
- Keep an eye out for drifting BPM positions
- Log all DAQ crashes in the excel worksheet along with the reason for the crash
- If beam delivery is interrupted for more than 4 hours, perform a checkout of the beam profile and raster follow Simona's procedure: 
- Cross check the rate with what is expected: 
- Cross check with the expected ion chamber readings: 
Notices in Effect for Shift Crew
- Power down SBS, BB, and Detector HV before controlled access and allowing RadCon in the hall
- Communicate to the TO that the SBS and BB magnets will be powered down, so they can update the holding field
- See the beam recovery procedures at: HOW_TOs#Beam_Recovery_Procedure
- See the instructions to set the raster at: HOW_TOs#How_to_ask_MCC_to_set_the_desired_raster_size_of_6_mm
- Shift Openings: Monday (morning) OWL, Tuesday OWL and SWING.